if you accelrate from rest to a rate of 6.00 m/s2 for a time of 4.10 seconds . determine displacement during the time period
1
Expert's answer
2012-10-04T08:25:10-0400
displacement is
s = (at^2)/2 = (6*4.1^2)/2 = 50.43 m